# Improved Approximation Algorithms for Chromatic and Pseudometric-Weighted Correlation Clustering

### v1

@@ 5.1 Pseudometric-weighted Correlation Clustering
$e.cost_{w}(u,v)$ =$=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-},\end{cases}$
$e.lp_{w}(u,v)$ =$=\begin{cases}(1-p_{uw}p_{vw})x_{uv},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}),&uv\in E^{-}.\end{cases}$
@@
Since the set of $(w_{uv},w_{vw},w_{wu})$ that satisfies the triangle inequality forms a convex cone generated by $(1,1,0),\,(1,0,1),\,(0,1,1)$, the function value is nonnegative for all such $(w_{uv},w_{vw},w_{wu})$ if and only if the value is nonnegative for $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$. ∎
This lemma implies that the algorithm achieves an approximation factor of $\alpha$ if all of the following inequalities are satisfied for every possible configuration on the triangle $uvw$:
@@
We obtain a lower bound on the approximation factor of LP-Pivot by verifying the feasibility of rounding functions that satisfy the above inequalities. To this end, we analyze several configurations of LP values and edge signs on triangle $uvw$.
@@
We begin by considering the configuration $(x,1-x,1)$ with signs (+,+,)(+,+,-)( + , + , - ) or (+,,)(+,-,-)( + , - , - ) and weights $(1,0,1)$. Then:
@@
Next, consider the configuration $(0,x,x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
@@
Now consider the configuration $(x,x,2x)$ with signs (+,+,+)(+,+,+)( + , + , + ) and weights $(1,1,0)$:
@@ 5.2 Chromatic Correlation Clustering
As in Section 5.1, the algorithm achieves an $\alpha$-approximation if the following inequality holds for all triangles $uvw$:
@@
Suppose LP-CCC is an expected $\alpha$-approximation algorithm, where $42\leq\alpha<$ 4.
Plugging $(\frac{1}{2},\frac{1}{2},1)$ with (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ), we obtain:
@@
Next, plugging $\left(\frac{1}{4},\frac{1}{4},\frac{1}{2}\right)$ with (+,+,)(+,+$\circ$)( + , + , ∘ ) yields:
@@
Using the known bound from chawla2015near , plugging $(0,x,x)$ with (+,+,+)(+,+,+)( + , + , + ) yields:
@@
On the other hand, from (24), the lower bound is $13-\alpha=3-4(1-t^{2})=4t^{2}-$ 1.
@@
Since $h^{\prime\prime}(t)=96t^{2}-66t+24>0$ and $h^{\prime}(0)=6>0$, $h$ is convex and increasing for $t\geq 0$.
@@ 5.3 Proof for Theorem 4
For sign configurations that do not involve a $\circ$’ (i.e., all edges have $+$ or -- signs), the inequality $\alpha\cdot LP-ALG\geq 0$ with $\alpha=2.06-\varepsilon$ is already established in chawla2015near . Thus, it suffices to check the remaining $6$ configurations involving at least one $\circ$’ sign.
@@
5.3.1 (,,)($\circ$,$\circ$,$\circ$)( ∘ , ∘ , ∘ ) Triangles
@@
$2\cdot LP-ALG\geq 0$
5.3.2 (,,+)($\circ$,$\circ$,+)( ∘ , ∘ , + ) Triangles
@@
5.3.3 (,,)($\circ$,$\circ$,-)( ∘ , ∘ , - ) Triangles
@@
Otherwise, since $21-z\leq 1/\alpha<1/$ 2,
@@
    Note that the function is affine with respect to $p_{y}$; therefore, the function is minimized when $p_{y}$ reaches the boundary, i.e., $y=1$ or $y=1-x$.
@@
    On the other hand, if $y=1-x$, then $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ as $y>1/2$, hence
@@
    Applying $y=1-x$, since $y=1-x>1/2$, $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ and thus
@@
The first term is a monotonically decreasing function of $x$; the second term is a monotonically decreasing function of $p_{x}$ since $-\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\leq 0$, as shown in the following Lemma:
@@
    is concave of $x$ since $\alpha\cdot 0.5095-2<0$ and $f^{+}$ is convex in the range. Hence, this is minimized at either $x=0$ or $x=0.5095$, where the former case results in 00 and the latter case results in $2\alpha\cdot 0.5095-2\geq 0$.
@@
    Since $z\geq 1/2$, $p_{z}>0.94>\frac{83}{89}$, thus the function is decreasing with respect to $p_{y}$. The minimum value of the function is therefore 00 when $p_{y}=p_{z}=1$.
@@
    We use the similar argument with of those conducted by Chawla et al.chawla2015near on the (+,+,)(+,+,-)( + , + , - ) case. For simplicity, let $a=0.19,\,b=0.5095$.
@@
    We now show that the function is minimized when $y=z$ as well as $x=y+z$. Consider the parametrization $(x,y,z)=(x,t-c,t+c)$, where $t\in[a,1/2],\,c\in[0,\min\{t-a,b-t\}],$ and $x=2t$ since $\alpha\cdot LP-ALG$ is a monotonically increasing function of $x$. Then
@@
    Fixing the value of $t$, the sum of first $2$ terms is a negative even quartic function of $c$, where the negativity is obtained from $-\alpha(1-t)+1\leq-\alpha(1-b)+1<0$. Thus, we can write the function as
@@
    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z=(\alpha(p_{y}-y)+2)z-(\alpha+1)p_{y}$
@@
    $-\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z\leq-\alpha y-p_{y}+2\leq-\frac{\alpha}{2}-f^{+}\left(\frac{1}{2}\right)+2\approx-0.016<0.$
@@
    $-\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\leq-\alpha+2<0.$
@@
    As mentioned above, it is sufficient to check on $x=1$ and $x=1-y$.
@@
      $x=1-y$
      Since $x=1-y>1/2$, $p_{x}=\frac{3}{10}x+\frac{7}{10}=1-\frac{3}{10}y$. Therefore,
@@
    since $z<1/2$ and $-2\alpha+2p_{x}+p_{y}\leq-2\alpha+3=-1.3<0$. Therefore, the function is minimized when $z$ is maximized.
@@ 6 Conclusion
In this work, we studied two important variants of correlation clustering: pseudometric-weighted correlation clustering and chromatic correlation clustering (CCC). For both problems, we developed and analyzed specialized rounding functions that are essential for achieving improved approximation guarantees via the LP-Pivot algorithm.
For the pseudometric-weighted setting, we proposed a piecewise-linear rounding function tailored for the setting that achieves a $10/3$-approximation, and showed that no alternative function within our analytical framework can yield a better factor. For the chromatic correlation clustering (CCC) variant, we designed a distinct rounding function that respects the constraints imposed by color restrictions and achieves an approximation factor of $2.15$. The function is constructed using a piecewise-linear form and leverages a careful analysis of triple costs.
@@
Appendix A Proof for $6$-approximation on Pseudometric-weighted CC with LP-UMVD-Pivot
Since the formulas for $e.cost$ and $e.lp$ in the pseudometric-weighted CC are compatible with those in the classical CC, and weights $w$ affect $\mathcal{C}$ linearly, the following lemma presented by Chawla et al. [16] remains applicable.
Lemma 5 (Lemma 5 of [16]).
  1. 1.
    the triangle inequality is tight for $(x_{uv},x_{vw},x_{wu})$; or
  2. 2.
We introduce another useful lemma:
Lemma 6.
Proof.
$\alpha\cdot e.lp_{w}-e.cost_{w}$ $\geq\alpha(1-p_{vw}p_{wu})\min\{x,1-x\}$
$\quad-\max\{p_{vw}(1-p_{wu})+(1-p_{vw})p_{wu},(1-p_{vw})(1-p_{wu})\}$
$\geq\alpha(1-p_{vw}p_{wu})\cdot\frac{1}{\alpha}-(1-p_{vw}p_{wu})=0.$
A.1 Triangle inequality is tight.
Figure 3: Configuration of regions induced from (34).
Table 1: Formula for $e.cost$ and $e.lp$ for each configuration.
$x_{uv}=x$ $x_{vw}=y$ $x_{wu}=z=x+y$
Region Sign $e.cost_{w}$ $e.lp_{w}$ $e.cost_{u}$ $e.lp_{u}$ $e.cost_{v}$ $e.lp_{v}$
I $+$ 00 $x$ 00 $y$ 00 $z$
-- $1$ $1-x$ $1$ $1-y$ $1$ $1-z$
II $+$ $z$ $x$ $z$ $y$ 00 $z$
-- $1-z$ $1-x$ $1-z$ $1-y$ $1$ $1-z$
III $+$ $y+z-2yz$ $(1-yz)x$ $z$ $y$ $y$ $z$
-- $(1-y)(1-z)$ $(1-yz)(1-x)$ $1-z$ $1-y$ $1-y$ $1-z$
IV $+$ $1-y$ $(1-y)x$ $1$ $y$ $y$ $z$
-- 00 $(1-y)(1-x)$ 00 $1-y$ $1-y$ $1-z$
V $+$ 00 00 $1$ $y$ $1$ $z$
-- 00 00 00 $1-y$ 00 $1-z$
VI $+$ $1-y$ $(1-y)x$ $1-x$ $(1-x)y$ $x+y-2xy$ $(1-xy)z$
-- 00 $(1-y)(1-x)$ 00 $(1-x)(1-y)$ $(1-x)(1-y)$ $(1-xy)(1-z)$
  1. 1.
  2. 2.
  3. 3.
  4. 4.
A.1.1 Region I, II, V
I $(0,0)$ $+$ 00 00 00
-- $5$ $5$ $5$
$(0,\frac{1}{3})$ $+$ 00 $2$ $2$
-- $5$ $3$ $3$
$(\frac{1}{6},\frac{1}{6})$ $+$ $1$ $1$ $2$
-- $4$ $4$ $3$
II $(0,\frac{1}{3})$ $+$ $-1/3$ $5/3$ $2$
-- $16/3$ $10/3$ $3$
$(\frac{1}{6},\frac{1}{6})$ $+$ $2/3$ $2/3$ $2$
-- $13/3$ $1/3$ $3$
$(\frac{1}{3},\frac{1}{3})$ $+$ $4/3$ $4/3$ $4$
-- $11/3$ $11/3$ $1$
V $(0,\frac{2}{3})$ $+$ 00 $3$ $3$
-- 00 $2$ $2$
$(0,1)$ $+$ 00 $5$ $5$
-- 00 00 00
$(\frac{1}{3},\frac{2}{3})$ $+$ 00 $3$ $5$
-- 00 $2$ 00
A.1.2 Region III
$y+z-2yz-(1-y)(1-z)=\frac{1-(2-3y)(2-3z)}{3}\geq 0$
since $y,z\in[1/3,2/3]$ within the region. Therefore, $6\cdot e.lp_{w}-e.cost_{w}$ is smaller with the sign ‘$+$’.
Now, consider the minimum value possible for the formula above within the region. Replacing $x=z-y$,
$6\cdot e.lp_{w}-e.cost_{w}=-6yz^{2}+6y^{2}z+2yz-7y+5z.$
Fixing $y$, the formula is concave of $z$; therefore, it is minimized in $z=y$ or $z=2/3$, resulting in $2y^{2}-2y$ and $4y^{2}-\frac{25}{3}y+\frac{10}{3}$ each. Within the range $y\in[1/3,2/3]$, the minimum value of each is $-1/2$ in $y=1/2$ and $-4/9$ in $y=2/3$. Thus, $26\cdot e.lp_{w}-e.cost_{w}\geq-1/$ 2.
Finally, consider the minimum value possible for $6\cdot e.lp_{u}-e.cost_{u}$ and $6\cdot e.lp_{v}-e.cost_{v}$. Again, these are affine; hence checking on extremal points is sufficient. It turns out that $36\cdot e.lp_{u}-e.cost_{u}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘$+$’ sign; $36\cdot e.lp_{v}-e.cost_{v}\geq 4/$ 3, minimized at $(y,z)=(1/3,2/3)$ with ‘--’ sign.
A.1.3 Region IV
We can utilize the optimization method explained in A.1.2 again: Since $(1-y)x\leq(1-y)(1-x)$ and $1-y\geq 0$, the formula with the sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$. This has a minimum value of $-1/3$ in $(x,y)=(0,2/3)$.
Also, since $26z-y=6x+5y\geq 10/3>5/$ 2 in this region, the formula with the sign ‘--’ is smaller for $6\cdot e.lp_{v}-e.cost_{v}$, which is $5-6x-5y$. This has a minimum value of $-1/3$ in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $16y-$ 1 or $6(1-y)$. Each of them has a minimum value $1$ in $(x,y)=(1/3,1/3)$, $2$ in $(x,y)=(0,2/3)$, respectively. Thus, $(6\cdot e.lp_{w}-e.cost_{w})+(6\cdot e.lp_{u}-e.cost_{u})$ and $(6\cdot e.lp_{u}-e.cost_{u})+(6\cdot e.lp_{v}-e.cost_{v})$ are at least $-1/3+1=2/3\geq 0$.
Finally,
$=-6xy-4y+4\geq-6y+4\geq 0,$
A.1.4 Region VI
As in IV, the formula with sign ‘$+$’ is smaller for $6\cdot e.lp_{w}-e.cost_{w}$, which is $(6x-1)(1-y)$. This has a minimum value of $1/3$ in $(x,y)=(1/3,2/3)$.
The formula for $6\cdot e.lp_{u}-e.cost_{u}$ is $(1-x)(6y-1)$ or $6(1-x)(1-y)$. Each of them has a minimum value $2/3$ in $(x,y)=(1/3,1/3)$, $4/3$ in $(x,y)=(1/3,2/3)$, respectively.
For $6\cdot e.lp_{v}-e.cost_{v}$, consider the difference between those of ‘$+$’ sign and ‘--’ sign.
$=xy-2x-2y+3$
$=(2-x)(2-y)-1\geq\frac{11}{9}>0,$
since $z\geq 2/3$. Therefore, the formula with sign ‘--’ is smaller, which is $6(1-xy)(1-z)-(1-x)(1-y)$.
Define $p=xy$, then
$6(1-xy)(1-z)-(1-x)(1-y)=6pz-5z-7p+5,$
with region $z\in[2/3,1],\,p\in\left[\frac{z}{3}-\frac{1}{9},\left(\frac{z}{2}\right)^{2}\right]$. Fixing the value of $z$, it is a decreasing function with respect to $p$; fixing the value of $p$, it is a decreasing function with respect to $z$. Therefore, the function is minimized at $z=1,\,p=\frac{1}{4}$ with the value $-\frac{1}{4}$.
Overall, $6\cdot LP-ALG\geq 0$ when the triangle inequality is tight.
A.2 All coordinates belong to endpoint of segments.
There are total $4$ possible LP values: 0, 1/3, 2/3, 10132310,\,1/3,\,2/3,\,10 , 1 / 3 , 2 / 3 , 1. Among all the possible $(x,y,z)$ tuples, there are only $7$ tuples whose triangle inequality is not tight. Moreover, $3$ tuples $(x,y,z)=(\frac{1}{3},\frac{1}{3},\frac{1}{3}),\,(\frac{1}{3},\frac{2}{3},\frac{2}{3}),\,(\frac{2}{3},\frac{2}{3},\frac{2}{3})$ among them only consist of value $1/3$ or $2/3$, thus $3\cdot LP-ALG\geq 0$ by Lemma 6. Therefore, we only need to verify the remaining $3$ tuples.
Since left-side and right-side limits of the value of rounding function at such point might differ, we have to consider all possible limits of $6\cdot LP-ALG$ up to $82^{3}=$ 8 directions. Table 3 shows remaining $3$ tuples and the directional limit of the value of $6\cdot e.cost-e.lp$ for each configuration. Note that since $x\leq y\leq z$, only up to $4$ out of $8$ directions are required to verification.
$(\frac{2}{3}-\delta,\frac{2}{3}-\delta,1)$ $+$ $1$ $1$ $26/9$
-- $2/3$ $2/3$ $-1/9$
$(\frac{2}{3}-\delta,\frac{2}{3}+\delta,1)$ $+$ 00 $1$ $5/3$
-- 00 $2/3$ 00
$(\frac{2}{3}+\delta,\frac{2}{3}+\delta,1)$ $+$ 00 00 00
-- 00 00 00
$(\frac{1}{3}-\delta,1,1)$ $+$ 00 $5$ $5$
-- 00 00 00
$(\frac{1}{3}+\delta,1,1)$ $+$ 00 $10/3$ $10/3$
-- 00 00 00
$(\frac{2}{3}-\delta,1,1)$ $+$ 00 $5/3$ $5/3$
-- 00 00 00
$(\frac{2}{3}+\delta,1,1)$ $+$ 00 00 00
-- 00 00 00
$(1,1,1)$ $+$ 00 00 00
-- 00 00 00
This completes our case analysis.
Appendix B Proof for Theorem 2
We can apply the same argument with Appendix A, with the partition by regions (Figure 4) and formula for $e.cost$ and $e.lp$ (Table 4) changed accordingly.

### v2

@@ 5.1 Pseudometric-weighted Correlation Clustering
$e.cost_{w}(u,v)$ $=\begin{cases}p_{uw}(1-p_{vw})+(1-p_{uw})p_{vw},&uv\in E^{+};\\ (1-p_{uw})(1-p_{vw}),&uv\in E^{-},\end{cases}$
$e.lp_{w}(u,v)$ $=\begin{cases}(1-p_{uw}p_{vw})x_{uv},&uv\in E^{+};\\ (1-p_{uw}p_{vw})(1-x_{uv}),&uv\in E^{-}.\end{cases}$
@@
Since the set of $(w_{uv},w_{vw},w_{wu})$ that satisfies the triangle inequality forms a convex cone generated by (1,1,0),(1,0,1),(0,1,1)(1,1,0),\,(1,0,1),\,(0,1,1), the function value is nonnegative for all such $(w_{uv},w_{vw},w_{wu})$ if and only if the value is nonnegative for $(w_{uv},w_{vw},w_{wu})\in\{(1,1,0),\,(1,0,1),\,(0,1,1)\}$. ∎
This lemma implies that the algorithm achieves an approximation factor of $\alpha$ if all of the following inequalities are satisfied for every possible configuration on the triangle uvwuvw:
@@
We obtain a lower bound on the approximation factor of LP-Pivot by verifying the feasibility of rounding functions that satisfy the above inequalities. To this end, we analyze several configurations of LP values and edge signs on triangle uvwuvw.
@@
We begin by considering the configuration (x,1x,1)(x,1-x,1) with signs (+,+,)(+,+,-) or (+,,)(+,-,-) and weights $(1,0,1)$. Then:
@@
Next, consider the configuration $(0,x,x)$ with signs $(+,+,+)$ and weights $(1,1,0)$:
@@
Now consider the configuration $(x,x,2x)$ with signs $(+,+,+)$ and weights $(1,1,0)$:
@@ 5.2 Chromatic Correlation Clustering
As in Section 5.1, the algorithm achieves an $\alpha$-approximation if the following inequality holds for all triangles uvwuvw:
@@
Suppose LP-CCC is an expected $\alpha$-approximation algorithm, where $2\leq\alpha<4$.
Plugging $(\frac{1}{2},\frac{1}{2},1)$ with $(\circ,\circ,-)$, we obtain:
@@
Next, plugging $\left(\frac{1}{4},\frac{1}{4},\frac{1}{2}\right)$ with $(+,+,\circ)$ yields:
@@
Using the known bound from chawla2015near , plugging $(0,x,x)$ with $(+,+,+)$ yields:
@@
On the other hand, from (25), the lower bound is $3-\alpha=3-4(1-t^{2})=4t^{2}-1$.
@@
Since $h^{\prime\prime}(t)=96t^{2}-66t+24>0$ and $^{\prime}(0)=6>0, hh$ is convex and increasing for $t\geq 0$.
@@ 5.3 Proof for Theorem 4
For sign configurations that do not involve a $\circ$’ (i.e., all edges have ++ or - signs), the inequality $\alpha\cdot LP-ALG\geq 0$ with $\alpha=2.06-\varepsilon$ is already established in chawla2015near . Thus, it suffices to check the remaining $6$ configurations involving at least one $\circ$’ sign.
@@
5.3.1 $(\circ,\circ,\circ)$ Triangles
@@
$2\cdot LP-ALG\geq 0.$
5.3.2 $(\circ,\circ,+)$ Triangles
@@
5.3.3 $(\circ,\circ,-)$ Triangles
@@
Otherwise, since $1-z\leq 1/\alpha<1/2$,
@@
    Note that the function is affine with respect to $p_{y}$; therefore, the function is minimized when $p_{y}$ reaches the boundary, i.e., $y=1$ or y=1xy=1-x.
@@
    On the other hand, if y=1xy=1-x, then $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ as $y>1/2$, hence
@@
    Applying y=1xy=1-x, since y=1x>1/2y=1-x>1/2, $p_{y}=\frac{3}{10}y+\frac{7}{10}=1-\frac{3}{10}x$ and thus
@@ 5.3.4 $(\circ,+,+)$ Triangles
The first term is a monotonically decreasing function of $x$; the second term is a monotonically decreasing function of $p_{x}$ since $\alpha(yp_{z}+p_{y}z)-2+2(p_{y}+p_{z})\leq 0$, as shown in the following Lemma:
@@
    is concave of $x$ since $\alpha\cdot 0.5095-2<0$ and $f^{+}$ is convex in the range. Hence, this is minimized at either $x=0$ or $x=0.5095$, where the former case results in 0 and the latter case results in $2\alpha\cdot 0.5095-2\geq 0$.
@@
    Since $z\geq 1/2$, $p_{z}>0.94>\frac{83}{89}$, thus the function is decreasing with respect to $p_{y}$. The minimum value of the function is therefore 0 when $p_{y}=p_{z}=1$.
@@
    We use the similar argument with of those conducted by Chawla et al.chawla2015near on the (+,+,)(+,+,-) case. For simplicity, let a=0.19,b=0.5095a=0.19,\,b=0.5095.
@@
    We now show that the function is minimized when $y=z$ as well as $x=y+z$. Consider the parametrization (x,y,z)=(x,tc,t+c)(x,y,z)=(x,t-c,t+c), where $t\in[a,1/2],\,c\in[0,\min\{t-a,b-t\}],$ and $x=2t$ since $\alpha\cdot LP-ALG$ is a monotonically increasing function of $x$. Then
@@
    Fixing the value of $t$, the sum of first $2$ terms is a negative even quartic function of $c$, where the negativity is obtained from $\alpha(1-t)+1\leq-\alpha(1-b)+1<0$. Thus, we can write the function as
@@ 5.3.5 $(\circ,+,-)$ Triangles
    $\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z=(\alpha(p_{y}-y)+2)z-(\alpha+1)p_{y}$
@@
    $\alpha yz-\alpha p_{y}(1-z)-p_{y}+2z\leq-\alpha y-p_{y}+2\leq-\frac{\alpha}{2}-f^{+}\left(\frac{1}{2}\right)+2\approx-0.016<0.$
@@
    $\alpha+1-\alpha y+\left(\frac{\alpha}{2}+1\right)p_{y}\leq-\alpha+2<0.$
@@
    As mentioned above, it is sufficient to check on $x=1$ and x=1yx=1-y.
@@
      x=1yx=1-y
      Since x=1y>1/2x=1-y>1/2, $p_{x}=\frac{3}{10}x+\frac{7}{10}=1-\frac{3}{10}y$. Therefore,
@@
    since $z<1/2$ and $2\alpha+2p_{x}+p_{y}\leq-2\alpha+3=-1.3<0$. Therefore, the function is minimized when $z$ is maximized.
@@ 6 Conclusion
In this work, we studied two important variants of correlation clustering: pseudometric-weighted correlation clustering and chromatic correlation clustering . For both problems, we developed and analyzed specialized rounding functions that are essential for achieving improved approximation guarantees via the LP-Pivot algorithm.
For the pseudometric-weighted setting, we proposed a piecewise-linear rounding function tailored for the setting that achieves a $10/3$-approximation, and showed that no alternative function within our analytical framework can yield a better factor. For the chromatic correlation clustering variant, we designed a distinct rounding function that respects the constraints imposed by color restrictions and achieves an approximation factor of $2.15$. The function is constructed using a piecewise-linear form and leverages a careful analysis of triple costs.
@@
Appendix A Proof for Theorem 2
We can apply the same argument with Appendix 4.4, with the partition by regions (Figure 4) and formula for $e.cost$ and $e.lp$ (Table 4) changed accordingly.